# A positive period derivative in the quasi-periodic eruptions of ZTF19acnskyy

### v1

@@ I Introduction
The leading model class thus far explains the recurring X-ray bursts via repeated interactions with a gravitationally captured stellar-mass object around the SMBH in an extreme mass-ratio inspiral (EMRI; e.g. King 2020; Xian et al. 2021; Suková et al. 2021; Krolik and Linial 2022; Linial and Sari 2023; Lu and Quataert 2023; Franchini et al. 2023; Linial and Metzger 2023; Tagawa and Haiman 2023; Yao et al. 2025)—though we will soon see the more appropriate term in our case may be extreme mass-ratio outspiral. In this picture, the bursts may arise from collisions of the orbiting companion with the accretion disk surrounding the SMBH (perhaps initially formed by the precursor TDE), or by repeated mass-loss of the companion near pericenter. EMRI models have provided an attractive framework to interpret the zeroth-order observational properties of QPEs, such as their amplitudes, quasi-periodicity, rates, and spectra. Still, there remain multiple open questions related to e.g. the precise source of short- and long-term timing variations observed in some QPEs (Miniutti et al., 2025; Arcodia et al., in prep.) and the burst energetics across the population (Mummery, 2025; Linial et al., 2025). Alternative models of accretion disk instabilities driven by e.g. nodal precession (Raj and Nixon, 2021; Middleton et al., 2025) or variations in magnetic viscosity (Śniegowska et al., 2020, 2023; Pan et al., 2022, 2023, 2025; Kaur et al., 2023) are thus worthwhile avenues for ongoing investigation. It is entirely possible that the growing family of QPEs may arise from a combination of different physical mechanisms.
The most extreme known source of QPEs is the nuclear transient ZTF19acnskyy/“Ansky” (Sánchez-Sáez et al., 2024), which hosts long, luminous eruptions of integrated energy budget $L\Delta t\gtrsim 100\times$ larger and recurrence time $\gtrsim 10\times$ longer than typical QPE sources (Hernández-García et al., 2025b). The eruptions also uniquely show broad, time-evolving absorption lines potentially indicating relativistic outflows powering the QPEs (Chakraborty et al., 2025c). The optical/UV transient nature of Ansky is uncertain, with possible interpretations including an AGN turn-on around a $\sim 10^{6}M_{\odot}$ SMBH or a featureless TDE (Sánchez-Sáez et al., 2024; Zhu et al., 2025). Furthermore, Hernández-García et al. (2025a) recently reported a doubled recurrence time and burst duration between 2024-2025, followed by a steady increase of $\sim 0.1$ d per flare. Ansky’s extraordinary characteristics provide a stress-test for the very limits of any physical model of QPEs.
@@ III Results
The bursts reach $L_{\rm peak}\approx(2-6)\times 10^{43}$ erg s-1, corresponding to time-integrated energy outputs $L\Delta t\approx(1-5)\times 10^{48}$ erg. Assuming radiative efficiencies in the range $\eta\sim 0.01-1$ (e.g. Linial et al. 2025), we find characteristic mass budgets of $\Delta M_{*}\sim L\Delta t/(\eta c^{2})\sim 10^{-6}-10^{-4}M_{\odot}$. Given that $\mathcal{O}(100)$ bursts are expected to have occurred thus far—including those observed in 2024 and 2026, and interpolating between observational gaps—we may infer a total mass budget on the order of $\sim 10^{-4}-10^{-2}M_{\odot}$ to date. In other words, the observed QPE properties and lifetime remain consistent with being powered by a stellar-mass object.
@@
where $T_{\mathrm{M},N}$ is the model timing of the $n$th epoch of model $N$; $\dot{P}$ is the period time-derivative; $\Delta P_{0}$ is a correction to the trial period $P_{0}$; and $A_{l/s}$, $P_{l/s}$, and $\phi_{l/s}$ are the amplitude, period, and phase of the long/short-period sinusoidal components in models 2-4. All four models and their residuals are plotted in Fig. 3.
Models 1 and 3 involve a series expansion of the QPE period about $P_{0}$, carried out to quadratic order (a constant $\dot{P}$). Models 2 and 4 consist only of sinusoidal modulation(s) of the burst arrival times, as is seen increasingly commonly in QPE timing analyses (Chakraborty et al., 2024a; Miniutti et al., 2025; Arcodia et al., in prep.). While the absolute values of $\chi^{2}$ are meaningless given our ad hoc choice of a $0.1$ d systematic error, the difference in reduced chi-squared can be used for qualitative model comparison. Models 1/2/3/4 have $\chi^{2}/\mathrm{dof.}=14.9/15.9/2.9/3.1$ respectively. Models 3 and 4 are thus favored over 1 and 2, while model 3—a period derivative plus a short-term sinusoidal modulation—achieves the best fit statistic.
@@ IV Discussion
Given the ambiguity between a constant $\dot{P}\approx 1.7\times 10^{-2}$ d d-1 or a long-term period oscillation over 112711-27 yr, as well as the variety of physical processes which may be relevant to period evolution in repeating nuclear transients (Linial and Quataert, 2024), we consider several models for the observed period change. In Sections IV.1-IV.2 we assume QPEs are powered by repeatedly stripping mass from a bound star near pericenter (akin to a repeating TDE, though not necessarily accretion-powered), which results in a steady increase of the orbital period as in models 1 & 3. In Sections IV.3-IV.4 we assume there is no secular change to the orbital period, but that the apparent $\dot{P}$ is the result of geometric light travel-time delays induced by general relativistic precession or the effect of a third body, à la models 2 & 4. Section IV.5 discusses the ability of disk instability models to describe the data.
@@ IV.1 Orbital evolution due to mass-transfer
Mass transfer from a lower-mass donor to a larger-mass companion can counteract angular momentum-loss mechanisms such as gravitational-wave emission or tidal dissipation and cause binary orbits to expand, an effect which has even been known theoretically for decades (see e.g. Eggleton 2006) and even measured directly in a few ultracompact binary systems via eclipse timing (de Miguel et al., 2018; Chakraborty et al., 2024b). This result has been extended to eccentric mass-transfering stellar binaries by deriving equations for the time evolution of the osculating orbital elements, e.g. via the orbit-averaged Lagrange planetary equations under a perturbing force due to mass-transfer (Hadjidemetriou, 1963; Sepinsky et al., 2009; Dosopoulou and Kalogera, 2016a, b). While the general case depends on the orbital angular momentum retention factor, size of the donor’s Roche lobe, donor/accretor masses, and other uncertain system properties, most of these are linear-order in the mass ratio $q\equiv M_{*}/M_{\bullet}$, and can thus be dropped due to the extreme mass ratio in our system ($q\sim 10^{-6}$). We follow Eqns. 43-44 of Dosopoulou and Kalogera (2016b), which assume delta-function Roche lobe overflow at pericenter and non-conservative mass transfer ($\dot{M}_{\rm tot},\dot{J}_{\rm orb}\neq 0$). In the EMRI approximation, the fractional change in orbital period ($P_{\rm orb}$) is then:
@@
where $e$ is the orbital eccentricity; $\Delta M_{*}$ is the fraction of the stellar mass loss per passage; and $\zeta$ is the fraction of stripped mass captured by the primary, defined such that $\Delta M_{\bullet}=-\zeta\Delta M_{*}$, as opposed to being lost to infinity and carrying orbital angular momentum away with it (e.g. as in Linial and Sari 2017). We note that Eq. 5 is reminiscent of the Newtonian result of Hughes 2019 obtained by enforcing adiabatic invariance of the orbital actions assuming conservative mass-transfer. Though the methods differ, they found $\Delta P_{\rm orb}/P_{\rm orb}=3(\Delta M_{*}/M_{*})$; our result differs by the order-unity factor $\sqrt{1-e^{2}}$ and the mass-retention term $\zeta$.
@@
implying the eccentricity is excited to larger values as the semimajor axis increases. In Fig. 5 we plot the evolution of $P_{\rm orb}$, $M_{*}$, and $e$ according to Eqns. 5-6 for various choices of ($e_{0}$, $\zeta$). The typical required fractional mass-loss per orbit is of order $\Delta M_{*}/M_{*}\approx(4-8)\times 10^{-3}$, increasing for lower initial eccentricities (as evident from Eq. 5). For a star of $\sim M_{\odot}$, this is a factor 10100010-1000 larger than estimated in Section II assuming a radiative efficiency of $\eta\sim 0.01-1$, but remarkably, is comparable to values found from the energetics of the time-evolving outflow reported in Ansky (Chakraborty et al., 2025c) implying $\eta$ may be lower than previously assumed. The inferred fractional mass loss carries the dramatic implication that $\sim 10\%$ of the stellar mass must be stripped within just a year, placing a strict bound on the continued eruption lifetime and predicting dramatic luminosity evolution should be seen in continued monitoring.
@@
[…]
@@ IV.3 General relativistic precession
Given the relatively constant burst energetics through 2025, it is worth considering whether the observed $\dot{P}$ may be caused by an oscillatory change to the system geometry and/or light travel time due to relativistic precession(s), rather than a true change to the orbital semimajor axis. These are expected to induce measurable effects on QPE timings (e.g. Xian et al. 2021; Linial and Metzger 2023, 2024; Franchini et al. 2023; Zhou et al. 2024), and such variations have already been observed in some cases (Chakraborty et al., 2024a; Miniutti et al., 2025; Arcodia et al., in prep.). Models 2 and 4 of Section III found a long-term precession period of 4001050400-1050 epochs (112711-27 yr) and an amplitude of 100045001000-4500 days; we consider whether the values can be reproduced.
@@
Further super-orbital variations, with longer timescales and larger amplitude, can occur due to Lense-Thirring (nodal) precession of the orbiter and/or the accretion disk (e.g. Franchini et al. 2023). While this is more difficult to model analytically, the effect can be readily simulated with the open-source EMRI/disk trajectory modeling code QPE-FIT [^1] https://github.com/joheenc/QPE-FIT (Chakraborty et al., 2025a), which performs Bayesian inference using QPE timings to infer EMRI/disk parameters while including relativistic precession of the orbiter and disk. We show the best-fitting models in Fig. 8: though the modulation amplitude can be approximately reproduced, the long modulation timescale cannot be captured simultaneously. This is the same tension we saw in Eq. 15: large-amplitude geometric delays require highly eccentric orbits, but larger $e$ shortens the precession timescales compared to the year baseline over which we observe a stable $\dot{P}$.
@@
It is worth noting the short-period oscillation in OCO-C models 3-4, with $P_{s}\approx 155$ d and amplitude $0.8$ d, is pluasibly consistent with relativistic precession. A 155 d period requires $e\gtrsim 0.97$ (Eq. 11)—which is not impossible (e.g. the orbit of S2 around Sgr A*)—and the amplitude $\ll T_{\rm rec}$ is significantly easier to obtain.
In any case, from the above arguments and the OCO-C constraints it appears unlikely that the apparent $\dot{P}$ is entirely the result of relativistic precession, due to the long modulation timescales (112711-27 yr) and amplitudes (100045001000-4500 days). It is possible that a large SMBH mass $\gg 10^{6}M_{\bullet}$ may increase the variation amplitude (e.g. Eq. 15), but this would both decrease the precession period (Eq. 11) and be in tension with host galaxy scaling relation-based estimates of $M_{\bullet}$ (Sánchez-Sáez et al., 2024; Zhu et al., 2025). It may thus be that reproducing the data requires a previously unconsidered source of precession-induced timing variation—an increasingly common conclusion of QPE timing studies (Miniutti et al., 2025; Arcodia et al., in prep.). While the most direct confirmation of this model would be eventually measuring the change to $\dot{P}$ predicted by long-term precession after a few years (Fig. 5), this may be achieved earlier by directly measuring $\ddot{P}$.

### v2

@@ I Introduction
The leading model class thus far explains the recurring X-ray bursts via repeated interactions with a gravitationally captured stellar-mass object around the SMBH in an extreme mass-ratio inspiral (EMRI; e.g. King 2020; Xian et al. 2021; Suková et al. 2021; Krolik and Linial 2022; Linial and Sari 2023; Lu and Quataert 2023; Franchini et al. 2023; Linial and Metzger 2023; Tagawa and Haiman 2023; Yao et al. 2025)—though we will soon see the more appropriate term in our case may be extreme mass-ratio outspiral. In this picture, the bursts may arise from collisions of the orbiting companion with the accretion disk surrounding the SMBH (perhaps initially formed by the precursor TDE), or by repeated mass-loss of the companion near pericenter. EMRI models have provided an attractive framework to interpret the zeroth-order observational properties of QPEs, such as their amplitudes, quasi-periodicity, rates, and spectra. Still, there remain multiple open questions related to e.g. the precise source of short- and long-term timing variations observed in some QPEs (Miniutti et al., 2025; Arcodia et al., subm.) and the burst energetics across the population (Mummery, 2025; Linial et al., 2025). Alternative models of accretion disk instabilities driven by e.g. nodal precession (Raj and Nixon, 2021; Middleton et al., 2025) or variations in magnetic viscosity (Śniegowska et al., 2020, 2023; Pan et al., 2022, 2023, 2025; Kaur et al., 2023) are thus worthwhile avenues for ongoing investigation. It is entirely possible that the growing family of QPEs may arise from a combination of different physical mechanisms.
The most extreme known source of QPEs is the nuclear transient ZTF19acnskyy/“Ansky” (Sánchez-Sáez et al., 2024), which hosts long, luminous eruptions of integrated energy budget $L\Delta t\gtrsim 100\times$ larger and recurrence time $\gtrsim 10\times$ longer than typical QPE sources (Hernández-García et al., 2025b). The eruptions also uniquely show broad, time-evolving absorption lines potentially indicating they are powered by relativistic outflows (Chakraborty et al., 2025c). The nature of the optical/UV transient preceding the X-ray QPEs in Ansky is uncertain, with possible interpretations including an AGN turn-on around a $\sim 10^{6}M_{\odot}$ SMBH or a featureless TDE (Sánchez-Sáez et al., 2024; Zhu et al., 2025). Ansky is the only QPE currently known to show a UV counterpart to the X-ray bursts (Guo et al., 2026), possibly as a result of the long timescales and large energy budget associated with the flares (Vurm et al., 2024; Chakraborty et al., 2025c). Furthermore, Hernández-García et al. (2025a) recently reported a doubled recurrence time and burst duration between 2024-2025, followed by a steady increase of $\sim 0.1$ d per flare. Ansky’s extraordinary characteristics provide a stress-test for the very limits of any physical model of QPEs.
@@ III Results
The bursts reach $L_{\rm peak}\approx(2-6)\times 10^{43}$ erg s-1, corresponding to time-integrated energy outputs $L\Delta t\approx(1-5)\times 10^{48}$ erg. Assuming radiative efficiencies in the range $\eta\sim 0.01-1$, we find characteristic mass budgets of $\Delta M_{*}\sim L\Delta t/(\eta c^{2})\sim 10^{-6}-10^{-4}M_{\odot}$ for accretion-powered emission, or $\Delta M_{*}\sim 10^{-3}-10^{-1}M_{\odot}$ for emission powered by shocks at velocities comparable to Keplerian for a circular orbit ($v_{\rm K}\sim 0.03c$). Given that $\mathcal{O}(100)$ bursts are expected to have occurred thus far—including those observed in 2024 and 2026, and interpolating between observational gaps—we may infer a total mass budget on the order of $10^{-4}-10^{-2}M_{\odot}$ for accretion-powered QPEs, or $0.1-10M_{\odot}$ for collisional shock-powered QPEs. The former model thus remains consistent with being powered by a stellar-mass object, while the latter may strain the total mass budget available within the stellar EMRI model (though see Linial et al. 2025).
@@
where $T_{\mathrm{M},N}$ is the model timing of the $n$th epoch of model $N$; $\dot{P}$ is the period time-derivative; $\Delta P_{0}$ is a correction to the trial period $P_{0}$; and $A_{l/s}$, $P_{l/s}$, and $\phi_{l/s}$ are the amplitude, period, and phase of the long/short-period sinusoidal components in models 2-4. All four models and their residuals are plotted in Fig. 3, with best-fitting parameters given in Table 2.
Models 1 and 3 involve a series expansion of the QPE period about $P_{0}$, carried out to quadratic order (a constant $\dot{P}$). Models 2 and 4 consist only of sinusoidal modulation(s) of the burst arrival times, as is seen increasingly commonly in QPE timing analyses (Chakraborty et al., 2024a; Miniutti et al., 2025; Arcodia et al., subm.). We note that for $n\ll P_{l}/P_{0}$, a Taylor expansion of the long-period sinusoidal component in models 2 and 4 yields a quadratic term $\propto n^{2}$, making models 1 and 2 (and 3 and 4) mathematically identical at quadratic order; the models become clearly distinguishable only after $\gtrsim 100$ further epochs (Fig. 4).
While the absolute values of $\chi^{2}$ are meaningless given our ad hoc choice of a $0.1$ d systematic error, the difference in reduced chi-squared can be used for qualitative model comparison. Models 1/2/3/4 have $\chi^{2}/\mathrm{dof.}=14.9/15.9/2.9/3.1$ respectively. Models 3 and 4 are thus favored over 1 and 2, while model 3—a period derivative plus a short-term sinusoidal modulation—achieves the best fit statistic.
Table 2: Best-fit parameters for the four OCO-C timing models. Dashes indicate parameters not present in a given model.
Parameter Model 1 Model 2 Model 3 Model 4
$T_{0}$ (MJD) $60710.02\pm 0.06$ $60710.09^{+0.09}_{-0.10}$
$P_{0}$ (d) $9.50\pm 0.01$ $9.48^{+0.02}_{-0.02}$
$\dot{P}$ (d d-1) $1.67^{+0.01}_{-0.01}\times 10^{-2}$ $1.72^{+0.01}_{-0.02}\times 10^{-2}$
$A_{l}$ (d) $1872^{+2309}_{-1077}$ $3068^{+1498}_{-1205}$
$P_{l}/P_{0}$ (cyc) $681^{+338}_{-238}$ $863^{+190}_{-190}$
$\phi_{l}$ (rad) $4.62^{+0.03}_{-0.04}$ $4.65^{+0.01}_{-0.02}$
$A_{s}$ (d) $0.79\pm 0.05$ $0.79\pm 0.05$
$P_{s}/P_{0}$ (cyc) $16.2^{+0.5}_{-0.5}$ $16.3^{+0.6}_{-0.6}$
$\phi_{s}$ (rad) $0.91^{+0.15}_{-0.13}$ $0.91\pm 0.15$
$\chi^{2}$/dof 297.7/20 302.1/19 49.3/17 49.6/16
$\chi^{2}_{\nu}$ 14.9 15.9 2.9 3.1
@@ IV Discussion
Given the formal degeneracy between a constant $\dot{P}\approx 1.7\times 10^{-2}$ d d-1 and the leading-order Taylor expansion of a long-period oscillation over 112711-27 yr with the current $\sim 30$-epoch baseline, as well as the variety of physical processes which may be relevant to period evolution in repeating nuclear transients (Linial and Quataert, 2024), we consider several models for the observed period change. In Sections IV.1-IV.2 we assume QPEs are powered by repeatedly stripping mass from a bound star near pericenter (akin to a repeating TDE, though not necessarily accretion-powered), which results in a steady increase of the orbital period as in models 1 & 3. In Sections IV.3-IV.4 we assume there is no secular change to the orbital period, but that the apparent $\dot{P}$ is the result of geometric light travel-time delays induced by general relativistic precession or the effect of a third body, à la models 2 & 4. Section IV.5 discusses the ability of disk instability models to describe the data.
@@ IV.1 Orbital evolution due to mass-transfer
Mass transfer from a lower-mass donor to a larger-mass companion can counteract angular momentum-loss mechanisms such as gravitational-wave emission or tidal dissipation and cause binary orbits to expand, an effect which has even been known theoretically for decades (see e.g. Eggleton 2006) and even measured directly in a few ultracompact binary systems via eclipse timing (de Miguel et al., 2018; Chakraborty et al., 2024b). This result has been extended to eccentric mass-transfering stellar binaries by deriving equations for the time evolution of the osculating orbital elements, e.g. via the orbit-averaged Lagrange planetary equations under a perturbing force due to mass-transfer (Hadjidemetriou, 1963; Sepinsky et al., 2009; Dosopoulou and Kalogera, 2016a, b). While the general case depends on the orbital angular momentum retention factor, size of the donor’s Roche lobe, donor/accretor masses, and other uncertain system properties, most of these are linear-order in the mass ratio $q\equiv M_{*}/M_{\bullet}$, and can thus be dropped due to the extreme mass ratio in our system ($q\sim 10^{-6}$). We follow Eqns. 43-44 of Dosopoulou and Kalogera (2016b), which assume delta-function Roche lobe overflow at pericenter and non-conservative mass transfer ($\dot{M}_{\rm tot},\dot{J}_{\rm orb}\neq 0$). We note that their formalism assumes mass is lost only through the inner Lagrange point L1, as is appropriate for stellar-mass binaries. In the extreme mass ratio limit ($q\sim 10^{-6}$), L2 mass loss becomes nontrivial (Linial and Sari 2017; Section IV.2), which will alter the angular momentum transfer and thus the magnitude of $\Delta P_{\rm orb}/P_{\rm orb}$ by an uncertain degree. Still, as an order-of-magnitude estimate, we can take the $q\ll 1$ limit to find the fractional change in orbital period ($P_{\rm orb}$):
@@
where $e$ is the orbital eccentricity; $\Delta M_{*}$ is the fraction of the stellar mass loss per passage; and $\zeta$ is the fraction of stripped mass captured by the primary, defined such that $\Delta M_{\bullet}=-\zeta\Delta M_{*}$, as opposed to being lost to infinity and carrying orbital angular momentum away with it (e.g. as in Linial and Sari 2017). We note that Eq. 5 is reminiscent of the Newtonian result of Hughes (2019) obtained by enforcing adiabatic invariance of the orbital actions assuming conservative mass-transfer. Though the methods differ, they found $\Delta P_{\rm orb}/P_{\rm orb}=3(\Delta M_{*}/M_{*})$; our result differs by the order-unity factor $\sqrt{1-e^{2}}$ and the mass-retention term $\zeta$.
@@
implying the eccentricity is excited to larger values as the semimajor axis increases. In Fig. 5 we plot the evolution of $P_{\rm orb}$, $M_{*}$, and $e$ according to Eqns. 5-6 for various choices of ($e_{0}$, $\zeta$). The typical required fractional mass-loss per orbit is of order $\Delta M_{*}/M_{*}\approx(4-8)\times 10^{-3}$, increasing for lower initial eccentricities (as evident from Eq. 5); remarkably, this mass budget is comparable to values inferred from the energetics of the time-evolving outflow reported in Ansky (Chakraborty et al., 2025c). The inferred fractional mass loss carries the dramatic implication that $\sim 10\%$ of the stellar mass must be stripped within just a year, placing a strict bound on the continued eruption lifetime and predicting dramatic luminosity evolution should be seen in continued monitoring.
@@
In comparison, the 2025 eruptions have durations of just 3 days. Reconciling the small required radii $\lesssim 14R_{g}$ (for $\alpha=0.1$ and $H/R=0.1$) with the orbital period and $M_{\bullet}$ would require an eccentricity of $e\approx 0.99$, precluding any stellar companion but a white dwarf (e.g. King 2020). Alternatively, a separate mechanism, such as collisional shocks of the stripped mass with the ambient disk, may be relevant for the system (e.g. Lu and Quataert 2023).
@@ IV.3 General relativistic precession
Given the relatively constant burst energetics through 2025, it is worth considering whether the observed $\dot{P}$ may be caused by an oscillatory change to the system geometry and/or light travel time due to relativistic precession(s), rather than a true change to the orbital semimajor axis. These are expected to induce measurable effects on QPE timings (e.g. Xian et al. 2021; Linial and Metzger 2023, 2024; Franchini et al. 2023; Zhou et al. 2024), and such variations have already been observed in some cases (Chakraborty et al., 2024a; Miniutti et al., 2025; Arcodia et al., subm.). Models 2 and 4 of Section III found a long-term precession period of 4001050400-1050 epochs (112711-27 yr) and an amplitude of 100045001000-4500 days; we consider whether the values can be reproduced.
@@
Further super-orbital variations, with longer timescales and larger amplitude, can occur due to Lense-Thirring (nodal) precession of the orbiter and/or the accretion disk (e.g. Franchini et al. 2023). While this is more difficult to model analytically, the effect can be readily simulated with the open-source EMRI/disk trajectory modeling code QPE-FIT [^1] https://github.com/joheenc/QPE-FIT v0.1.11 (Chakraborty et al., 2025a), which performs Bayesian inference using QPE timings to infer EMRI/disk parameters while including relativistic precession of the orbiter and disk. We show the best-fitting models in Fig. 8: though the modulation amplitude can be approximately reproduced, the long modulation timescale cannot be captured simultaneously. This is the same tension we saw in Eq. 15: large-amplitude geometric delays require highly eccentric orbits, but larger $e$ shortens the precession timescales compared to the year baseline over which we observe a stable $\dot{P}$.
@@
It is worth noting the short-period oscillation in OCO-C models 3-4, with $P_{s}\approx 155$ d and amplitude $0.8$ d, is plausibly consistent with relativistic precession. A 155 d period requires $e\gtrsim 0.97$ (Eq. 11)—which is not impossible (e.g. the orbits of S-stars around Sgr A*, Burkert et al. 2024)—and the amplitude $\ll T_{\rm rec}$ is significantly easier to obtain.
In any case, from the above arguments and the OCO-C constraints it appears unlikely that the apparent $\dot{P}$ is entirely the result of relativistic precession, due to the long modulation timescales (112711-27 yr) and amplitudes (100045001000-4500 days). It is possible that a large SMBH mass $\gg 10^{6}M_{\bullet}$ may increase the variation amplitude (e.g. Eq. 15), but this would both decrease the precession period (Eq. 11) and be in tension with host galaxy scaling relation-based estimates of $M_{\bullet}$ (Sánchez-Sáez et al., 2024; Zhu et al., 2025). It may thus be that reproducing the data requires a previously unconsidered source of precession-induced timing variation—an increasingly common conclusion of QPE timing studies (Miniutti et al., 2025; Arcodia et al., subm.). While the most direct confirmation of this model would be eventually measuring the change to $\dot{P}$ predicted by long-term precession after a few years (Fig. 5), this may be achieved earlier by directly measuring $\ddot{P}$.